OpenStudy (anonymous):

A bag contains 2 white marbles and 7 purple marbles. Two marbles are drawn at random. One marble is drawn and not replaced. Then a second marble is drawn. 1.) what is the probability of selecting a purple marble and then a white marble? 2.)what is the probability of selecting two white marbles? 3.) Is there a greater chance of selecting two white marbles in a row or two purple marbles in a row?

OpenStudy (gebooors):

There are 9 marbles. Question 1: p(first purple) = 7/9 then there are 8 marbles left. P(second is white) = 2/8. Then multiply probabilities. 7/9 * 2/8 = 14/72 = 7/36 Try to solve next questions in a same way.

OpenStudy (academicgurusinc):

P(White)= Probability of selecting white on 1st draw is 7/13 P(Purple)= Probability of selecting purple on 2nd draw is 6/12. The total marble is now 12 because you do not replace the 1st marble. the probability that the first marble is white and the second one is purple is P(White)*(P(Purple). Hope that helps. If you found this helpful, I encourage to subscribe to our youtube channel https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg to stay current with all of our new videos. Regards, Academic Gurus Inc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)